When ABC was trading at $52 per share, you paid $4.20 for a put option (for one share) on the stock of ABC with a strike price of $50, and six months until maturity. After six months, the share price of ABC is $54.10. A) What is the value of the put option at expiration? Do not include the $ sign and answer to the nearest $0.01. $ B) What is your profit or loss? Losses should be entered as negative numbers. Do not include the $ sign and answer to the nearest $0.01.
Answer #1

A) Put option gives owner right to sell asset at predetermined price in given time frame

The strike price = $ 50

Value of put option = Max(Strike price - spot price at expiry, 0 ) = Max (50 -54.10 , 0) = 0.00

B)  Profit/(Loss)= Value of put option - premium

Profit/(Loss)= 0.00 - 4.20 = (4.20)